# Edwards-Wilkinson limit for a stochastic advection-diffusion PDE

### v1

@@ 2.1. Correlation Functions and a priori estimates
Proof.
Following [10], we denote $\theta^{[M]}(t,x)$ as the solution of the stochastic transport-diffusion equation (1.1) starting at t=Mt=-M with initial data $\theta^{[M]}(-M,\cdot)\equiv 1$; then it holds (cf. [10, Corollary 3.2])
@@
Note that by construction, the stationary field $\Psi(n^{2}(r-s),x)$ depends on information of the driving field $V(t,x)$ for $t<n^{2}(r-s)<0$, while $\theta^{[n^{2}(r-s)]}(0,x)$ depends on the information of $V(t,x)$ for $t\geqslant n^{2}(r-s)$; thus, they are independent. As a result, the second expectation is equal to
@@
Here we prove Theorem 1.9, under Assumption 1.1; in this case, $V_{\rm eff}^{2}=\int_{\mathbb{R}^{d}}Q(x)\,\mathrm{d}x$.
@@
(2.26) $\langle\tilde{M}_{n},\phi\rangle]_{t}=\int_{0}^{t}\big{\langle}\tilde{\theta}^{n}_{s}\nabla\phi,Q_{n}\ast(\tilde{\theta}^{n}_{s}\nabla\phi)\big{\rangle}\,\mathrm{d}s.$
We wish to show that
(2.27) $\langle\tilde{M}_{n},\phi\rangle]_{t}\rightarrow\int_{[0,t]}\int_{\mathbb{R}^{d}}\big{(}q_{r}\ast\varphi(x)\big{)}^{2}\nabla\phi(x)^{T}V_{\rm eff}^{2}\nabla\phi(x)\,\mathrm{d}x\,\mathrm{d}r,$
in probability, where now, $V_{\rm eff}^{2}=\int_{\mathbb{R}^{d}}Q(x)\,\mathrm{d}x$. We split the right-hand side of (2.26) as follows:
@@
$\leqslant\|\varphi\|_{2}\|\nabla\phi\|_{\infty}\int_{0}^{t}\big{\|}Q_{n}\ast(\bar{\theta}_{s}\nabla\phi)-V_{\rm eff}^{2}(\bar{\theta}_{s}\nabla\phi)\big{\|}_{2}\,\mathrm{d}s$
which vanishes as $n\to\infty$ since $Q_{n}=n^{d}Q(n\cdot)$ is similar to an approximation of identity on $\mathbb{R}^{d}$, with $\int_{\mathbb{R}^{d}}Q(x)\,\mathrm{d}x=V_{\rm eff}^{2}$.
@@
which also vanishes as $n\to\infty$. In the same way, we can show that $J_{3}\to 0$.
@@
By Proposition 2.6, for some $\varepsilon\in(0,1)$, we have
@@
$\tilde{\mathbb{E}}|J_{1}|$ $\leqslant\tilde{\mathbb{E}}\int_{0}^{T}\big{\|}(\tilde{\theta}^{n}_{s}-\bar{\theta}_{s})\nabla\phi\big{\|}_{2}\big{\|}Q_{n}\ast((\tilde{\theta}^{n}_{s}-\bar{\theta}_{s})\nabla\phi)\big{\|}_{2}\,\mathrm{d}s$
$\leqslant\tilde{\mathbb{E}}\int_{0}^{T}\big{\|}(\tilde{\theta}^{n}_{s}-\bar{\theta}_{s})\nabla\phi\big{\|}_{2}^{2}\|Q_{n}\|_{1}\,\mathrm{d}s,$
where in the second step we have used Young’s inequality. Next, since $\|Q_{n}\|_{1}=\|Q\|_{1}$ and $\tilde{\theta}^{n}$ has the same law as $\theta^{n}$, we have
$\tilde{\mathbb{E}}|J_{1}|$ $\leqslant\|\nabla\phi\|_{\infty}^{2}\|Q\|_{1}\,\tilde{\mathbb{E}}\int_{0}^{T}\big{\|}\tilde{\theta}^{n}_{s}-\bar{\theta}_{s}\big{\|}_{2}^{2}\,\mathrm{d}s$
$=\|\nabla\phi\|_{\infty}^{2}\|Q\|_{1}\,\mathbb{E}\int_{0}^{T}\big{\|}\theta^{n}_{s}-\bar{\theta}_{s}\big{\|}_{2}^{2}\,\mathrm{d}s\to 0$
as $n\to\infty$.
@@
Acknowledgements. The second author is grateful to the financial supports of the National Key R&D Program of China (No. 2024YFA1012300) and the National Natural Science Foundation of China (Nos. 12090010, 12090014).

### v2

@@ 2.1. Correlation Functions and a priori estimates
Remark 2.15.
In the following proof, we need to work with certain solutions to (1.7) defined at negative times. For this, we introduce suitable filtrations. We extend the i.i.d. Brownian motions $B_{k}$ to i.i.d. two-sided Brownian motions, still denoted by $B_{k}$ (letting $B_{k}(t)=B^{\prime}_{k}(-t)$ for $t<0$, where $\{B^{\prime}_{k}\}_{k}$ is a family of i.i.d. Brownian motions independent on $\{B_{k}\}_{k}$). For $\infty<s\leqslant t<\infty$, we call $\mathcal{F}_{s,t}$ the $\sigma$-algebra generated by $\{B_{k}(r)-B_{k}(s)\}_{k\in\mathbb{N},r\in[s,t]}$ and the $\mathbb{P}$-null sets; we also call $\mathcal{G}_{t}$ the sigma-algebra generated by $\{\mathcal{F}_{s,t}\}_{s\leqslant t}$. Clearly $\mathcal{G}_{s}$ and $\mathcal{F}_{s,t}$ are independent for every $s<t$.
Following [17], we consider the solution $\theta^{[M]}$ to the transport-diffusion equation (1.7) starting at t=Mt=-M with initial data $\theta^{[M]}(-M,\cdot)\equiv 1$, more precisely $\theta^{[M]}$ is defined by
$\theta^{[M]}=\mathbb{E}\Big[\big(\det DX^{[M]}_{t}\big(X^{[M],-1}_{t}(x)\big)\big)^{-1}\,\big|\,\sigma(\mathcal{F}_{-M,r},r\geq-M)\Big]$
where $X^{[M]}_{t}$ is the stochastic flow solving
$dX^{[M]}_{t}(x)=\sum_{k}\sigma_{k}\big(X^{[M]}_{t}(x)\big)\circ\,\mathrm{d}B_{k}(t)+\sqrt{2\kappa}\,dW(t),\quad X^{[M]}_{-M}(x)=x,$
with $W$ an independent $d$-dimensional Brownian motion. Then $\theta^{[M]}_{t}$ is $(\mathcal{F}_{-M,t})$-measurable. Moreover, by (2.24) below, the stationary field $\Psi(t,\cdot)$ is $\mathcal{G}_{t}$-measurable.
Proof of Lemma 2.14.
Let $\theta^{[M]}(t,x)$ be the solution to (1.7) starting at t=Mt=-M with initial data $\theta^{[M]}(-M,\cdot)\equiv 1$, as in Remark 2.15. Then it holds (cf. [17, Corollary 3.2])
@@
Note that by construction, the stationary field $\Psi(n^{2}(r-s),x)$ depends on information of the driving field $V(t,\cdot)$ for $t<n^{2}(r-s)<0$ (precisely, it is $\mathcal{G}_{n^{2}(r-s)}$-measurable), while $\theta^{[n^{2}(r-s)]}(0,x)$ depends on the information of $V(t,x)$ for $t\geqslant n^{2}(r-s)$ (precisely, it is $\mathcal{F}_{n^{2}(r-s),0}$-measurable); thus, they are independent. As a result, the second expectation is equal to
@@
Here we prove Theorem 1.9, under Assumption 1.1; in this case, $V_{\rm eff}^{2}=g(0)\Pi$, where the function $g$ is given in (1.5) and $\Pi$ is the Helmholtz-Leray projection.
@@
(2.26) $\langle\tilde{M}_{n},\phi\rangle]_{t}=\int_{0}^{t}\big\langle\tilde{\theta}^{n}_{s}\nabla\phi,Q_{n}\ast(\tilde{\theta}^{n}_{s}\nabla\phi)\big\rangle\,\mathrm{d}s,$
where $Q_{n}(x)=n^{d}Q(nx)$, $x\in\mathbb{R}^{d}$. We wish to show that
(2.27) $\langle\tilde{M}_{n},\phi\rangle]_{t}\rightarrow\int_{0}^{t}\!\!\int_{\mathbb{R}^{d}}(\theta_{s}\nabla\phi)(x)^{\ast}V_{\rm eff}^{2}(\theta_{s}\nabla\phi)(x)\,\mathrm{d}x\,\mathrm{d}s,$
in probability, where now, $V_{\rm eff}^{2}=g(0)\Pi$.
We split the right-hand side of (2.26) as follows:
@@
$\leqslant\|\varphi\|_{2}\|\nabla\phi\|_{\infty}\int_{0}^{t}\big\|Q_{n}\ast(\bar{\theta}_{s}\nabla\phi)-V_{\rm eff}^{2}(\bar{\theta}_{s}\nabla\phi)\big\|_{2}\,\mathrm{d}s.$
By the Plancherel identity,
$\big\|Q_{n}\ast(\bar{\theta}_{s}\nabla\phi)-V_{\rm eff}^{2}(\bar{\theta}_{s}\nabla\phi)\big\|_{2}^{2}$ $=\int_{\mathbb{R}^{d}}\big|\widehat{Q}_{n}\mathcal{F}(\bar{\theta}_{s}\nabla\phi)-g(0)P_{\xi}\mathcal{F}(\bar{\theta}_{s}\nabla\phi)\big|^{2}\,\mathrm{d}\xi$
$=\int_{\mathbb{R}^{d}}\big|\big(g(n^{-1}\xi)-g(0)\big)P_{\xi}\mathcal{F}(\bar{\theta}_{s}\nabla\phi)\big|^{2}\,\mathrm{d}\xi,$
where we have used $\widehat{Q}_{n}(\xi)=\widehat{Q}(n^{-1}\xi)=g(n^{-1}\xi)P_{\xi}$ by (1.5). Recall that $g\in C_{b}(\mathbb{R}^{d})$; for any $s\in[0,t]$ and $\xi\neq 0$, it is clear that the integrand vanishes as $n\to\infty$, thus by the dominated convergence theorem, one has
$\lim_{n\to\infty}\big\|Q_{n}\ast(\bar{\theta}_{s}\nabla\phi)-V_{\rm eff}^{2}(\bar{\theta}_{s}\nabla\phi)\big\|_{2}=0.$
Therefore, the first term of $J_{4}$ vanishes as $n\to\infty$.
@@
which, similarly to the treatment of the first term in $J_{4}$, also vanishes as $n\to\infty$. In the same way, we can show that $J_{3}\to 0$.
@@
By Proposition 2.7 (and Remark 2.8), for some $\varepsilon\in(0,1)$, we have
@@
$\tilde{\mathbb{E}}|J_{1}|$ $\leqslant\tilde{\mathbb{E}}\int_{0}^{T}\big\|(\tilde{\theta}^{n}_{s}-\bar{\theta}_{s})\nabla\phi\big\|_{2}\big\|Q_{n}\ast((\tilde{\theta}^{n}_{s}-\bar{\theta}_{s})\nabla\phi)\big\|_{2}\,\mathrm{d}s;$
by Plancherel’s identity and (1.5),
$\big\|Q_{n}\ast((\tilde{\theta}^{n}_{s}-\bar{\theta}_{s})\nabla\phi)\big\|_{2}^{2}$ $=\int_{\mathbb{R}^{d}}\big|\widehat{Q}_{n}(\xi)\mathcal{F}((\tilde{\theta}^{n}_{s}-\bar{\theta}_{s})\nabla\phi)(\xi)\big|^{2}\,\mathrm{d}\xi$
$\leqslant\|g\|_{\infty}\int_{\mathbb{R}^{d}}\big|\mathcal{F}((\tilde{\theta}^{n}_{s}-\bar{\theta}_{s})\nabla\phi)(\xi)\big|^{2}\,\mathrm{d}\xi$
$=\|g\|_{\infty}\big\|(\tilde{\theta}^{n}_{s}-\bar{\theta}_{s})\nabla\phi\big\|_{2}^{2}.$
Plugging this estimate in the above inequality yields
$\tilde{\mathbb{E}}|J_{1}|$ $\leqslant\|g\|_{\infty}\tilde{\mathbb{E}}\!\int_{0}^{T}\!\big\|(\tilde{\theta}^{n}_{s}-\bar{\theta}_{s})\nabla\phi\big\|_{2}^{2}\,\mathrm{d}s\leqslant\|g\|_{\infty}\|\nabla\phi\|_{\infty}\tilde{\mathbb{E}}\!\int_{0}^{T}\!\big\|\tilde{\theta}^{n}_{s}-\bar{\theta}_{s}\big\|_{2}^{2}\,\mathrm{d}s.$
By Lemma 2.16 and the fact that $\tilde{\theta}^{n}$ has the same law as $\theta^{n}$, we conclude that the last quantity vanishes as $n\to\infty$.
@@
Acknowledgements. The first author acknowledges the partial support of the project Noise in fluid dynamics and related models, funded by the Italian Ministry of University and Research through the project PRIN call 2022 - grant 20222YRYSP. The second author is grateful to the financial supports of the National Key R&D Program of China (No. 2024YFA1012300) and the National Natural Science Foundation of China (Nos. 12090010, 12090014). The third author acknowledges the partial support of the project Convergence and Stability of Reaction and Interaction Network Dynamics, funded by the Italian Ministry of University and Research through the project PRIN call 2022 - grant 2022XRWY7W.